Session: Mar 2010 (session 1)
Wavelength: 18 cm


To give feedback on this experiment, click here

Experiment Feedback

Medicina (A.Maccaferri at Sat Mar 06 14:56:48 2010) Success:

Onsala (Michael Lindqvist at Sat Mar 06 18:21:45 2010) Success:
Apart from RFI, no known problems.

Noto (P. Cassaro at Sun Mar 07 11:10:22 2010) Failure: FS
Problems with the Field System

Westerbork (Geert Kuper at Sun Mar 07 13:12:46 2010) Success: RFI PHAS
12 RT's in the array. (No RT1 and RT5) Severe RFI on 1610.49 MHz both upper and lower sideband. This may have caused a bad phase tuneup for the first two sub bands

Effelsberg (dg at Mon Mar 08 09:58:54 2010) Minor failures: WIND
Lost most of experiment due to heavy snow. Start at 10:02. Gain may be affected by snow in antenna

Torun (Kaz Borkowski at Mon Mar 08 13:06:59 2010) Success:
No known problems.

Noto (P. Cassaro at Wed Mar 10 11:26:43 2010) Minor failures:
ADDENDUM: performed with the rubidium clock

Jodrell_Bank (alastair gunn at Mon Mar 29 10:50:14 2010) Success: ANT
Frequent telescope encoder faults giving brief periods off source. No other known problems.

Noto (P. Cassaro at Mon May 10 09:28:21 2010) Failure:
The schedule was not able to read the setup01 procedure, then the obseration was performed without a valid setup. No idea why.